Suppose Z^ - hZ1 A where Z1,A are independent random variables with mean 0 Part a) If σ-4, σ:1-1 and φ- 2? 0.3 , what is σ Part b) If σ 6 and φ-0.7, what is ơỈ in order that σ}i-Ž ?
Answer #1

Part a)

sigma^2_{Z_2} = Var[Z_2] = Var[phi Z_1 + A] = phi^2 Var[Z_1] + Var[A] (A, Z1 are independent random variables)

sigma^2_{Z_2} = phi^2 sigma^2_{Z_1}+ sigma^2_A = 0.3^2 * 1 + 4 = 4.09

Part b)

From part a),

sigma^2_{Z_2} = phi^2 sigma^2_{Z_1}+ sigma^2_A

For sigma^2_{Z_2} = sigma^2_{Z_1} ,

sigma^2_{Z_1} = phi^2 sigma^2_{Z_1}+ sigma^2_A

Rightarrow sigma^2_A = sigma^2_{Z_1} - phi^2 sigma^2_{Z_1}

Rightarrow sigma^2_A =(1 - phi^2) sigma^2_{Z_1}

= (1-0.72) * 6 = 3.06